# Shanghai 200433, China

### v1

@@ 1 Introduction
Symmetry is another important feature of physical systems, making it natural to study entanglement through block-diagonalization in the presence of symmetry as follows. Consider a bipartite system with global U(1) symmetry with conserved charge $Q$, which is a sum of charge of the two subsystems, $Q=Q_{A}+Q_{B}$. When the total system density matrix $\rho$ commutes with the conserved charge $Q$, it can be proved that, by partial trace, the subsystem density matrix still commute with the corresponding charge operator, i.e., $[\rho_{A},Q_{A}]=[\rho_{B},Q_{B}]=0$. Therefore, the subsystem density matrix $\rho_{A}$ will be block-diagonal in the eigenbases of the conserved charge $Q_{A}$:
@@
where each $\tilde{\rho}_{A}(q)$ is the projection of $\rho_{A}$ onto a charge sector. It is natural to study the entanglement entropy within each charge sector, called symmetry-resolved entanglement entropy (SREE) Goldstein:SRE
@@
$S(\rho_{A})=-\sum_{i}p(q)\log p_{A}(q)+\sum_{q}p(q)S(\rho_{A}(q)),$ (4)
The first term in this decomposition formula is the Shannon entropy of the probability distribution $\{p(q)\}$, and the second term is the average entanglement entropy of all sectors. For a more detailed explanation of symmetry resolution of entanglement and a perspective closer to Shannon’s original idea Shannon , readers may refer to Appendix A.
@@
$\operatorname{tr}[\tilde{\rho}_{A}(q)^{n}]\equiv\mathcal{Z}_{n}(q)=\int_{0}^{2\pi}\frac{\mathrm{d}\mu}{2\pi}\mathcal{Z}_{n}(\mu)\mathrm{e}^{-\mathrm{i}\mu q}.$ (8)
@@
In quantum field theory, $\mathcal{Z}_{n}(\mu)$ is computed as a path integral, with 2-dimensional conformal field theory (CFT) offering tractable examples Goldstein:SRE . Some other results of charged moments and SREE includes computations in Wess-Zumino-Witten (WZW) models Calabrese:SREE_in_WZW and in excited states of CFT Capizzi:SRE_of_excited , and analysis for the charged moment in 2-dimensional Dirac and complex scalar fields Murciano:SR_2DFreeQFT , among others. In the framework of AdS/CFT correspondence Maldacena:AdSCFT ; Witten:AdS_and_holography ; GKP , some proposals of the holographic dual of the charged moment has been presented in Zhao:SRE_in_AdSCFT , which was later examined in difference cases Weisenberger:SRE_excited_and_2interval . Besides entanglement entropy, other entanglement measures were also discussed in the context of symmetry-resolved entanglement, such as the symmetry-resolution of relative entropy Chen:SR_rela ; Capizzi:SR_rela , the capacity of entanglement Arias:capacity , and mutual information Parez:ExactQuench_SRE . For other related studies on symmetry-resolved entanglement, see Horvath:SRE_integrable ; Zhao:W3_higher_spin ; Ares:multi_chargedMoment ; Murciano:SR_Page ; DiGiulio:BCFT_to_SRE ; Berthiere:reflEntrp_CrsNormNegat ; Kusuki:SREE_BCFT ; Rottoli:Ising_crossover ; Pirmoradian:SRE_local_nonLocal ; Feldman:SRE_lattice ; Bianchi:nonAbel_SREE ; Choi:noninvert_SR_ALC .
@@
Besides advancements in exploring symmetry-resolved entanglement measures in (1+1)-dimensional cases, several studies have also discussed higher-dimensional cases, such as Fermi gas in 2 spatial dimensions Tan:particleNumFluct_2DFermiGas , free non-relativistic massless fermions and free bosons in 2 spatial dimensions by dimensional reduction Murciano:SRE_2D_dimReduct , and free Fermi gas in general dimensions using Widom conjecture Fraenkel:SRE_1D_beyond . However, most studies explored cases in specific dimensions, and the universal structures of symmetry-resolved entanglement in general dimensions is not clear. In this paper, we propose a method to compute SREE that can be applied in general higher dimensions. The method mainly relies on the well-known technique of computing entanglement entropy for a spherical region in a CFT by conformal mapping Myers:cthm_holo ; CHM:towards ; Hung:holo_renyi . The conformal mapping relates the domain of dependence of the spherical region to the hyperbolic cylinder ×$\mathbb{R}\times H^{d-1}$. To compute charged moment, we also employ the method in Belin:holo_charged_renyi , incorporating the new insertion $\mathrm{e}^{\mathrm{i}\mu Q}$ in the conformal mapping, and then the charged moment will be converted to a grand-canonical partition function on the hyperbolic space. While calculating this partition function is still difficult for generic CFT, we examine two types of theories where explicit computation is feasible. The first type is free field theories, in which computation of partition function, and hence the charged moment, can be done exactly, and then SREE can be computed from the charged moments following steps as stated above. The second type of theories in which the partition function can be computed is holographic theories. Following Belin:holo_charged_renyi  [^1] Strictly speaking, the quantities considered in Belin:holo_charged_renyi are not exactly the same as symmetry-resolved entanglement measures. However, the techniques can still apply here., the grand-canonical partition function will be equivalent to the grand-canonical partition function of a charged black hole due to $\mathrm{e}^{\mathrm{i}\mu Q}$ in the charged moment. Due to the universal applicability of the conformal mapping, our method can be applied to general dimensions.
@@
[…]
@@
The paper is organized as follows: In Section 2, we introduce the method of computing SREE of a spherical region in higher-dimensional CFTs. In Section 3, we apply this method to free field theory to compute SREE. In Section 4, we apply this method to holographic CFT in higher dimensions to compute SREE. The results in Section 3 and 4 give the same universal asymptotic structure of charged moments, SRRE, and SREE. In particular, the property of equipartition of SREE at leading orders is again discovered. In Section 5, we summarize and analyze the universal expansion structure of SRRE and SREE, and show that the structure is a result of choosing the size of the subsystem as the expansion parameter, and details of the underlying field theory is irrelevant. We conclude with a discussion in Section 6. In Appendix A, we motivate symmetry-resolved entanglement from the perspective of entropy coarse-graining, following the primitive thoughts of Shannon. In Appendix B, we provide the explanation and derivation of an asymptotic expansion formula used in explicit computations.
@@ 2 Methods
Consider a $d$-dimensional CFT in flat space in its vacuum state, and choose a spherical region $A$ of radius $R$ as a subsystem. Via the conformal mapping presented in CHM:towards , the domain of dependence $\mathcal{D}$ of the spherical region $A$ can be mapped to a hyperbolic cylinder ×$\mathbb{R}\times H^{d-1}$, where $\mathbb{R}$ represents the time direction and $H^{d-1}$ represents $(d-1)$-dimensional hyperbolic space. Furthermore, the reduced density matrix of the spherical region is mapped to the thermal state density matrix in the hyperbolic space, with temperature being $T_{0}=\beta_{0}^{-1}=\frac{1}{2\pi R}$:
@@
We also note that there is a simplification in the computation. As shown in Equation (9), the calculation of SRRE requires only the ratio $\frac{\mathcal{Z}_{n}(q)}{\mathcal{Z}_{1}^{n}(q)}$ rather than individual values of $\mathcal{Z}_{n}(q)$. It can be easily shown that, since the normalization factor $Z(\beta_{0},0)^{n}$ is a $\mu$-independent number raised to the power of $n$, it will cancel out in the ratio. Therefore, we can use unnormalized charged moments, defined as $\tilde{\mathcal{Z}}_{n}(\mu):=Z\left(n\beta_{0},\frac{\mathrm{i}\mu}{n\beta_{0}}\right)$, and correspondingly define the unnormalized symmetry-resolved partition function as
$\tilde{\mathcal{Z}}_{n}(q):=\int_{0}^{2\pi}\frac{\mathrm{d}\mu}{2\pi}\tilde{\mathcal{Z}}_{n}(\mu)\mathrm{e}^{-\mathrm{i}\mu q}=\int_{0}^{2\pi}\frac{\mathrm{d}\mu}{2\pi}Z\left(n\beta_{0},\frac{\mathrm{i}\mu}{n\beta_{0}}\right)\mathrm{e}^{-\mathrm{i}\mu q}.$ (19)
@@ 3 Free Field Computation
As mentioned in the last section, computing the charged moment $\mathcal{Z}_{n}(\mu)=\operatorname{tr}[\rho^{n}\mathrm{e}^{\mathrm{i}\mu Q}]$ is equivalent to computing the partition function with imaginary chemical potential. In the path integral approach, this imaginary chemical potential can be seen as modifying the boundary condition of the field Belin:holo_charged_renyi ; Roberge:gauge_Im . For example, a scalar boson traversing the Euclidean time circle will acquire an extra phase Roberge:gauge_Im ; Belin:holo_charged_renyi ; Goldstein:SRE :
@@
The modification to the boundary condition is captured effectively by the heat kernel method Belin:holo_charged_renyi , which also yields a phase $\mathrm{e}^{\mathrm{i}\mu}$ around the Euclidean time direction:
@@
where $x$ denotes the coordinate along the periodic Euclidean time direction $S_{\beta}^{1}$ (the subscript $\beta$ indicates the periodicity), and $y$ represents the coordinates on $H^{d-1}$. Note that $t$ in the above formula represent a parameter in the heat kernel, rather than time. For convenience, we now set the hyperbolic space radius to be dimensionless $R=1$. Consequently, the temperature relevant to $\mathcal{Z}_{n}(\mu)=\operatorname{tr}[\rho^{n}\mathrm{e}^{\mathrm{i}\mu Q}]$ is $T=\frac{1}{2\pi n}$.
@@
where the heat kernel on $S^{1}_{\beta}$ (which should satisfy the twisted boundary conditions (26)) is given by Grigor'yan:heatKer_hyperbolic ; Vassilevich:heatkernel_manual ; Belin:holo_charged_renyi
@@
Following the normalization discussion in Section 2, we calculate SREE using the unnormalized charged moment $\tilde{\mathcal{Z}}_{n}(\mu)=\mathrm{e}^{-F_{n}(\mu)}$, where free energy $F_{n}(\mu)$ can be expressed in terms of the heat kernel as
$F_{n}(\mu)=-\frac{1}{2}\int_{0}^{+\infty}\frac{\mathrm{d}t}{t}\int\mathrm{d}x\,\mathrm{d}^{d-1}y\,K_{{S^{1}_{\beta}\times H^{d-1}}}(\{x,y\},\{x,y\},t).$ (29)
In the following, we will consider specific theories to determine the explicit form of the heat kernel $K_{S^{1}_{\beta}\times H^{d-1}}$, which will allow us to compute the free energy and charged moment, and ultimately obtain the SREE.
@@
where $\mathcal{R}$ is the Ricci scalar curvature, and its coupling with $|\phi|^{2}$ with a coefficient $\frac{1}{6}$ ensures the conformal symmetry in 4D. We work with complex scalar rather than real scalar because complex scalar possesses global U(1) symmetry but real scalar does not possess such global symmetry. The heat kernel of the complex scalar field is twice as large as that of a real scalar field. The heat kernel for a real scalar field is given by the heat kernel of the Laplace operator (taking into account conformal coupling) on the corresponding manifold. For odd-dimensional hyperbolic space $H^{d-1}$, the expression of real scalar heat kernel is Grigor'yan:heatKer_hyperbolic
@@
The free energy of the scalar field on $S_{\beta}^{1}\times H^{3}$ is given by
[…]
@@
Several points about this expression need explanation. Firstly, the factor of $2$ in front of the heat kernel appears because a complex scalar field has twice the degrees of freedom of a real scalar field, as mentioned above. Secondly, the $k=0$ term in the sum produces a divergence proportional to $n$ and independent of $\mu$, and hence it will be eventually cancelled in the ratio $\frac{\mathcal{Z}_{n}(q)}{\mathcal{Z}_{1}^{n}(q)}$ when computing SRRE. Therefore, the $k=0$ term is omitted in the second line, and the remaining terms sum to polylogarithm functions $\mathrm{Li}_{n}(z)$. Thirdly, the hyperbolic space volume $V$ is also divergent; we formally retain it as the symbol $V$, which can later be assigned a finite value using cutoff regularization. Finally, the free energy $F_{n}(\mu)$ is periodic, and within the period $\mu\in[0,2\pi]$, the sum of the two polylogarithm functions simplifies as
@@
Next, we need to perform inverse Fourier transform (19) from $\mathcal{Z}_{n}(\mu)$ to $\mathcal{Z}_{n}(q)$. Since the transform generally lacks an analytical solution, we employ asymptotic analysis, using the volume of hyperbolic space $V$ as the expansion parameter. Performing asymptotic analysis (the details of this technique is explained in Appendix B), we obtain the unnormalized symmetry-resolved partition function to the subleading order
@@
This relationship can be interpreted as the equivalence of ensembles in the thermodynamic limit. While we are examining entanglement in a quantum system, the conformal mapping enables us to analyze it from a thermodynamic perspective. From this perspective, $S$ represents the total entropy of the system in a ensemble permitting charge fluctuation, whereas $S(q)$ is in an “micro-canonical” ensemble in which charge takes a certain value. In the thermodynamic limit, which corresponds to the large subsystem limit, charge fluctuation becomes negligible, and $S$ and $S(q)$ agree at the leading order. In formula, it follows from the definition of symmetry-resolved Rényi entropy in Equation (6):
@@
The subsequent terms represent the difference between symmetry-resolved entanglement entropy and conventional entanglement entropy, corresponding to the second term in Equation (42). Among these terms, the highest-order divergence is the logarithmic term $-\frac{1}{2}\log V$. The coefficient $-\frac{1}{2}$ comes from the asymptotic behavior of $F_{n}(\mu)$ at $\mu=0$. More specifically, the first derivative of $F_{n}(\mu)$ vanishes at $\mu=0$, while the second derivative does not. This causes $\sqrt{V}$ in the denominator in Equation (36) and finally $-\frac{1}{2}\log V$ in $S_{n}(q)$ and $S(q)$.
Motivated by the universal terms in entanglement entropy, we explore whether analogous universal contributions might arise in SREE. The logarithmic term $-\frac{1}{45}\log(R/\delta)$ in original entanglement entropy is inherently a universal term, whose coefficient arises from type A anomaly Solodukhin:EE_Conf_ExtGeom ; Fursaev:SquashCone ; and the leading order of $\log V$ also contributes a logarithmic term $-\frac{1}{2}\log\left(\frac{R^{d-2}}{\delta^{d-2}}\right)=-\log(R/\delta)$, in which $d=4$ in this case. These two logarithmic terms combine to give a term $S_{\log}(q)=-\frac{46}{45}\log\left(\frac{R}{\delta}\right)$, which is unaffected when changing $\delta$. Despite this robust mathematical structure, the physical significance of this combination of type A anomaly and dimensions remains unclear. Further work is needed to determine whether they encode meaningful universal data or reflect subtleties inherent to the SREE construction itself.
@@
In summary, for a conformal free complex scalar field in 4D, the expansion of symmetry-resolved entanglement entropy $S(q)$ in terms of $V$ consists of the conventional entanglement entropy $S$, plus the subleading term $-\frac{1}{2}\log V$, followed by further subleading terms. The equipartition property holds up to the $\mathcal{O}(V)$ term, and breaks down at the next higher-order term. Notably, the expansion structure of SREE aligns with the 2D compact scalar field results shown in Equation (10). In particular, the equipartition properties holds up to the constant order. In Section 5, we will show that the expansion structure is a consequence of the expansion in $V$, and that the dependence on $q$ can only appear in terms beyond the constant order.
@@
[…]
@@
As evident in Figure 2, $\mu=0$ is a minimum point of $F_{n}(\mu)$ and hence a maximum point of $\tilde{\mathcal{Z}}_{n}(\mu)$, and is thus consistent with the fact that $\left|\tilde{\mathcal{Z}}_{n}(\mu)/\tilde{\mathcal{Z}}_{n}(0)\right|<1$ for $\mu\neq 2k\pi,\,(k\in\mathbb{Z})$, as analyzed in 4D case.
@@
Among subsequent terms, the highest-order divergence is again the logarithmic term $-\log V$, but with a coefficient $-1$ different from $-\frac{1}{2}$ in 4D case. This results from the denominator $\sqrt{V}$ in $\tilde{\mathcal{Z}}_{n}(q)$, which is a result of non-vanishing $F_{n}^{\prime}(0)$, as mentioned above. The universal term in $S(q)$ arises only from the original entanglement entropy:

### v2

@@ 1 Introduction
Symmetry is another important feature of physical systems, making it natural to study entanglement through block-diagonalization in the presence of symmetry as follows. Consider a bipartite system with global U(1) symmetry with conserved charge $Q$, which is a sum of charge of the two subsystems, $Q=Q_{A}+Q_{B}$. When the total system density matrix $\rho$ commutes with the conserved charge $Q$, it can be proved that, by partial trace, the subsystem density matrix still commute with the corresponding charge operator, i.e., $\rho_{A},Q_{A}]=[\rho_{B},Q_{B}]=0$. Therefore, the subsystem density matrix $\rho_{A}$ will be block-diagonal in the eigenbases of the conserved charge $Q_{A}$:
@@
where each $\tilde{\rho}_{A}(q)$ is the projection of $\rho_{A}$ onto the sector with charge $q$, and the probability for the subsystem $A$ occupying the $q$ sector is given by $p_{A}(q)=\operatorname{tr}[\tilde{\rho}_{A}(q)]$. It is natural to study the entanglement entropy within each charge sector, called symmetry-resolved entanglement entropy (SREE) Goldstein:SRE
@@
$S(\rho_{A})=-\sum_{q}p_{A}(q)\log p_{A}(q)+\sum_{q}p_{A}(q)S(\rho_{A}(q)),$ (4)
The first term in this decomposition formula is the Shannon entropy of the probability distribution $\{p_{A}(q)\}$, and the second term is the average entanglement entropy of all sectors. For a more detailed explanation of symmetry resolution of entanglement and a perspective closer to Shannon’s original idea Shannon , readers may refer to Appendix A.
@@
$\operatorname{tr}[\tilde{\rho}_{A}(q)^{n}]\equiv\mathcal{Z}_{n}(q)=\int_{-\pi}^{\pi}\frac{\mathrm{d}\mu}{2\pi}\mathcal{Z}_{n}(\mu)\mathrm{e}^{-\mathrm{i}\mu q}.$ (8)
@@
In quantum field theory, $\mathcal{Z}_{n}(\mu)$ is computed as a path integral, with 2-dimensional conformal field theory (CFT) offering tractable examples Goldstein:SRE . Some other results of charged moments and SREE includes computations in Wess-Zumino-Witten (WZW) models Calabrese:SREE_in_WZW and in excited states of CFT Capizzi:SRE_of_excited , and analysis for the charged moment in 2-dimensional Dirac and complex scalar fields Murciano:SR_2DFreeQFT , among others. In the framework of AdS/CFT correspondence Maldacena:AdSCFT ; Witten:AdS_and_holography ; GKP , the holographic dual of the charged moment (initially proposed in Belin:holo_charged_renyi from a perspective distinct from SREE) has been explicitly linked to SREE in Zhao:SRE_in_AdSCFT . The proposal in Zhao:SRE_in_AdSCFT was later examined in Weisenberger:SRE_excited_and_2interval . Some studies have extended symmetry resolution of entanglement beyond conventional group-like symmetries to categorical symmetries Saura-Bastida:cat_SRE ; Das:gener_SRE ; Choi:noninvert_SR_ALC . Besides entanglement entropy, other entanglement measures were also discussed in the context of symmetry-resolved entanglement, such as the symmetry resolution of relative entropy Chen:SR_rela ; Capizzi:SR_rela , the capacity of entanglement Arias:capacity , and mutual information Parez:ExactQuench_SRE . For other related studies on symmetry-resolved entanglement, see Horvath:SRE_integrable ; Zhao:W3_higher_spin ; Baiguera:shape_deform_charged_Renyi ; Ares:multi_chargedMoment ; Murciano:SR_Page ; DiGiulio:BCFT_to_SRE ; Berthiere:reflEntrp_CrsNormNegat ; Kusuki:SREE_BCFT ; Rottoli:Ising_crossover ; Pirmoradian:SRE_local_nonLocal ; Feldman:SRE_lattice ; Bianchi:nonAbel_SREE .
@@
Besides advancements in exploring symmetry-resolved entanglement measures in (1+1)-dimensional cases, several studies have also discussed higher-dimensional cases, such as Fermi gas in 2 spatial dimensions Tan:particleNumFluct_2DFermiGas , free non-relativistic massless fermions and free bosons in 2 spatial dimensions by dimensional reduction Murciano:SRE_2D_dimReduct , and free Fermi gas in general dimensions using Widom conjecture Fraenkel:SRE_1D_beyond . A result for the charged moments across a hyperplane in $d$ Euclidean dimensions has also been presented Murciano:SR_2DFreeQFT . However, most studies explored cases in specific dimensions, and the universal structures of symmetry-resolved entanglement in general dimensions is not clear. In this paper, we propose a method to compute SREE that can be applied in general higher dimensions. The method mainly relies on the well-known technique of computing entanglement entropy for a spherical region in a CFT by conformal mapping Myers:cthm_holo ; CHM:towards ; Hung:holo_renyi . The conformal mapping relates the domain of dependence of the spherical region to the hyperbolic cylinder $\mathbb{R}\times H^{d-1}$. To compute charged moment, we also employ the method in Belin:holo_charged_renyi , incorporating the new insertion $\mathrm{e}^{\mathrm{i}\mu Q}$ in the conformal mapping, and then the charged moment will be converted to a grand-canonical partition function on the hyperbolic space. While calculating this partition function is still difficult for generic CFT, we examine two types of theories where explicit computation is feasible. The first type is free field theories, in which computation of partition function, and hence the charged moment, can be done exactly, and then SREE can be computed from the charged moments following steps as stated above. The second type of theories in which the partition function can be computed is holographic theories. Following Belin:holo_charged_renyi  [^1] Strictly speaking, the quantities considered in Belin:holo_charged_renyi are not exactly the same as symmetry-resolved entanglement measures. However, the techniques can still apply here., the grand-canonical partition function will be equivalent to the grand-canonical partition function of a charged black hole due to $\mathrm{e}^{\mathrm{i}\mu Q}$ in the charged moment. Due to the universal applicability of the conformal mapping, our method can be applied to general dimensions.
@@
$S(q)=S-\#\log V+\mathcal{O}\left(V^{0}\right)+q\text{-dependent terms},$ (11)
@@
The paper is organized as follows: In Section 2, we introduce the method of computing SREE of a spherical region in higher-dimensional CFTs. In Section 3, we apply this method to free field theory to compute SREE. In Section 4, we apply this method to holographic CFT in higher dimensions to compute SREE. The results in Section 3 and 4 give the same universal asymptotic structure of charged moments, SRRE, and SREE. In particular, the property of equipartition of SREE at leading orders is again discovered. In Section 5, we summarize and analyze the universal expansion structure of SRRE and SREE, and show that the structure is a result of choosing the size of the subsystem as the expansion parameter, and details of the underlying field theory is irrelevant. We conclude with a discussion in Section 6. In Appendix A, we motivate symmetry-resolved entanglement from the perspective of entropy coarse-graining, following the primitive thoughts of Shannon. In Appendix C, we provide the explanation and derivation of an asymptotic expansion formula used in explicit computations.
@@ 2 Methods
Consider a $d$-dimensional CFT in flat space in its vacuum state, and choose a spherical region $A$ of radius $R$ as a subsystem. Via the conformal mapping presented in CHM:towards , the domain of dependence $\mathcal{D}$ of the spherical region $A$ can be mapped to a hyperbolic cylinder $\mathbb{R}\times H^{d-1}$, where $\mathbb{R}$ represents the time direction and $H^{d-1}$ represents (d1)(d-1)-dimensional hyperbolic space. Furthermore, the reduced density matrix of the spherical region is mapped to the thermal state density matrix in the hyperbolic space, with temperature being $T_{0}=\beta_{0}^{-1}=\frac{1}{2\pi R}$:
@@
We also note that there is a simplification in the computation. As shown in Equation (9), the calculation of SRRE requires only the ratio $\frac{\mathcal{Z}_{n}(q)}{\mathcal{Z}_{1}^{n}(q)}$ rather than individual values of $\mathcal{Z}_{n}(q)$. It can be easily shown that, since the normalization factor $Z(\beta_{0},0)^{n}$ is a $\mu$-independent number raised to the power of $n$, it will cancel out in the ratio. Therefore, we can use unnormalized [^3] In this language, $\mathcal{Z}_{n}(\mu)$ is called “normalized” charged moment, since $\mathcal{Z}_{1}(0)=1$. charged moments, defined as $\tilde{\mathcal{Z}}_{n}(\mu):=Z\left(n\beta_{0},\frac{\mathrm{i}\mu}{n\beta_{0}}\right)$, and correspondingly define the unnormalized symmetry-resolved partition function as
$\tilde{\mathcal{Z}}_{n}(q):=\int_{-\pi}^{\pi}\frac{\mathrm{d}\mu}{2\pi}\tilde{\mathcal{Z}}_{n}(\mu)\mathrm{e}^{-\mathrm{i}\mu q}=\int_{-\pi}^{\pi}\frac{\mathrm{d}\mu}{2\pi}Z\left(n\beta_{0},\frac{\mathrm{i}\mu}{n\beta_{0}}\right)\mathrm{e}^{-\mathrm{i}\mu q}.$ (19)
@@ 3 Free Field Computation
As mentioned in the last section, computing the charged moment $\mathcal{Z}_{n}(\mu)=\operatorname{tr}[\rho^{n}\mathrm{e}^{\mathrm{i}\mu Q}]$ is equivalent to computing the partition function with imaginary chemical potential. For free field computations in this section, the partition function and the charged moment can be expressed in terms of the heat kernel Vassilevich:heatkernel_manual . Here we outline the ingredients needed for these computations. A pedagogical introduction to the heat kernel, along with related results, is provided in Appendix B.
In the path integral approach, this imaginary chemical potential can be seen as modifying the boundary condition of the field Belin:holo_charged_renyi ; Roberge:gauge_Im . For example, a scalar boson traversing the Euclidean time circle will acquire an extra phase Roberge:gauge_Im ; Belin:holo_charged_renyi ; Goldstein:SRE :
@@
The modification to the boundary condition is captured effectively by the heat kernel, which also yields a phase $\mathrm{e}^{\mathrm{i}\mu}$ around the Euclidean time direction Belin:holo_charged_renyi :
@@
where $x$ denotes the coordinate along the periodic Euclidean time direction $S_{\beta}^{1}$ (the subscript $\beta$ indicates the periodicity), and $y$ represents the coordinates on $H^{d-1}$. Note that $t$ in the above formula represents a parameter in the heat kernel, rather than time. For convenience, we now set the hyperbolic space radius to be dimensionless $R=1$. Consequently, the temperature relevant to $\mathcal{Z}_{n}(\mu)=\operatorname{tr}[\rho^{n}\mathrm{e}^{\mathrm{i}\mu Q}]$ is $T=\frac{1}{2\pi n}$.
@@
where the factor on $S^{1}_{\beta}$ (which should satisfy the twisted boundary conditions (26)) is given by Grigor'yan:heatKer_hyperbolic ; Vassilevich:heatkernel_manual ; Belin:holo_charged_renyi (see Appendix B for a derivation)
@@
Following the normalization discussion in Section 2, we calculate SREE using the unnormalized charged moment $\tilde{\mathcal{Z}}_{n}(\mu)=\mathrm{e}^{-F_{n}(\mu)}$. For complex scalar theories considered in the following subsections, the free energy $F_{n}(\mu)$ can be expressed in terms of the heat kernel as (for more details, see the following subsections)
$F_{n}(\mu)=-\int_{0}^{+\infty}\frac{\mathrm{d}t}{t}\int\mathrm{d}x\,\mathrm{d}^{d-1}y\,K_{{S^{1}_{\beta}\times H^{d-1}}}(\{x,y\},\{x,y\},t).$ (29)
We now consider specific theories to determine the explicit form of $K_{S^{1}_{\beta}\times H^{d-1}}$, which will allow us to compute the free energy and charged moment, and ultimately obtain the SREE.
@@
where $\mathcal{R}$ is the Ricci scalar curvature, and its coupling with $|\phi|^{2}$ with a coefficient $\frac{1}{6}$ ensures the conformal symmetry in 4D. We consider a complex scalar field rather than a real scalar field, since the former admits a global U(1) symmetry $\phi\rightarrow\phi\,\mathrm{e}^{\mathrm{i}\alpha}$ but the (non-compact) real scalar does not. The kinetic differential operator in this theory is the Laplace operator supplemented by a conformal coupling term. On an odd-dimensional hyperbolic space $H^{d-1}$, the corresponding heat kernel is Grigor'yan:heatKer_hyperbolic
@@
The heat kernel on $S_{\beta}^{1}$ is given by Equation (28), and we provide a derivation of it in Appendix B.
The free energy of the complex scalar field on $S_{\beta}^{1}\times H^{3}$ is given by
$F_{n}(\mu)$ $=-\int_{0}^{+\infty}\frac{\mathrm{d}t}{t}\int_{0}^{2\pi n}\mathrm{d}x\int\mathrm{d}^{3}y\,K_{{S^{1}_{\beta}\times H^{3}}}(\{x,y\},\{x,y\},t)$ (33)
@@
Several points about this expression need explanation. Firstly, the $k=0$ term in the sum produces a divergence proportional to $n$ and independent of $\mu$, and hence it will be eventually cancelled in the ratio $\frac{\mathcal{Z}_{n}(q)}{\mathcal{Z}_{1}^{n}(q)}$ when computing SRRE. Therefore, the $k=0$ term is omitted in the second line, and the remaining terms sum to polylogarithm functions $\mathrm{Li}_{n}(z)$. Secondly, the hyperbolic space volume $V$ is also divergent; we formally retain it as the symbol $V$, which can later be assigned a finite value using cutoff regularization. Finally, the free energy $F_{n}(\mu)$ is periodic, and within the period $\mu\in[0,2\pi]$, the sum of the two polylogarithm functions simplifies as
@@
Next, we need to perform the inverse Fourier transform (19) from $\mathcal{Z}_{n}(\mu)$ to $\mathcal{Z}_{n}(q)$. Since the transform generally lacks an analytical solution, we employ asymptotic analysis, using the volume of hyperbolic space $V$ as the expansion parameter. Performing asymptotic analysis (the details of this technique is explained in Appendix C), we obtain the unnormalized symmetry-resolved partition function to the subleading order
@@
This relationship can be interpreted as the equivalence of ensembles in the thermodynamic limit. While we are examining entanglement in a quantum system, the conformal mapping enables us to analyze it from a thermodynamic perspective. From this perspective, $S$ represents the total entropy of the system in an ensemble permitting charge fluctuation, whereas $S(q)$ is in a “micro-canonical” ensemble in which charge takes a certain value. In the thermodynamic limit, which corresponds to the large subsystem limit, charge fluctuation becomes negligible, and $S$ and $S(q)$ agree at the leading order. In formula, it follows from the definition of symmetry-resolved Rényi entropy in Equation (6):
@@
The subsequent terms represent the difference between symmetry-resolved entanglement entropy and conventional entanglement entropy, corresponding to the second term in Equation (42). Among these terms, the highest-order divergence is the logarithmic term $\frac{1}{2}\log V$. The coefficient $\frac{1}{2}$ comes from the asymptotic behavior of $F_{n}(\mu)$ at $\mu=0$. More specifically, the first derivative of $F_{n}(\mu)$ vanishes at $\mu=0$, while the second derivative does not. This causes $\sqrt{V}$ in the denominator in Equation (36) and finally $\frac{1}{2}\log V$ in $S_{n}(q)$ and $S(q)$.
Motivated by the universal terms in entanglement entropy, we explore whether analogous universal contributions might arise in SREE. The logarithmic term $\frac{1}{45}\log(R/\delta)$ in original entanglement entropy is inherently a universal term, whose coefficient arises from type A anomaly Solodukhin:EE_Conf_ExtGeom ; Fursaev:SquashCone ; and the leading order of $\log V$ also contributes a logarithmic term $\frac{1}{2}\log\left(\frac{R^{d-2}}{\delta^{d-2}}\right)=-\log(R/\delta)$, in which $d=4$ in this case. These two logarithmic terms combine to give a term $S_{\log}(q)=-\frac{46}{45}\log\left(\frac{R}{\delta}\right)$, which is unaffected when changing $\delta$. Despite this robust mathematical structure, the physical significance of this combination of type A anomaly and dimensions remains unclear. Further work is needed to determine whether they encode meaningful universal data or reflect subtleties inherent to the SREE construction itself.
@@
In summary, for a conformal free complex scalar field in 4D, the expansion of symmetry-resolved entanglement entropy $S(q)$ in terms of $V$ consists of the conventional entanglement entropy $S$, plus the subleading term $\frac{1}{2}\log V$, followed by further subleading terms. The equipartition property holds up to the $\mathcal{O}(V)$ term, and breaks down at the next higher-order term. Notably, the expansion structure of SREE aligns with the 2D compact scalar field results shown in Equation (10). In particular, the equipartition properties holds up to the constant order. In Section 5, we will show that the expansion structure is a consequence of the expansion in $V$, and that the dependence on $q$ can only appear in terms beyond the constant order.
@@
$F_{n}(\mu)$ $=-\int_{0}^{+\infty}\frac{\mathrm{d}t}{t}\int_{0}^{2\pi n}\mathrm{d}x\int\mathrm{d}y\,K_{S_{\beta}^{1}\times H^{1}}\left(\{x,y\},\{x,y\},t\right)$ (44)
@@
As evident in Figure 2, $\mu=0$ is a minimum point of $F_{n}(\mu)$ and hence a maximum point of $\tilde{\mathcal{Z}}_{n}(\mu)$, and is thus consistent with the fact that $\left|\tilde{\mathcal{Z}}_{n}(\mu)/\tilde{\mathcal{Z}}_{n}(0)\right|<1$ for $\mu\neq 2k\pi\,(k\in\mathbb{Z})$, as analyzed in 4D case.
@@
Among subsequent terms, the highest-order divergence is again the logarithmic term $\log V$, but with a coefficient 1-1 different from $\frac{1}{2}$ in 4D case. This results from the denominator $\sqrt{V}$ in $\tilde{\mathcal{Z}}_{n}(q)$, which is a result of non-vanishing $F_{n}^{\prime}(0)$, as mentioned above. The universal term in $S(q)$ arises only from the original entanglement entropy: